# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ 4.1. $L^{2}$-uniform bound of the logarithmic derivative $\mathsf{d}^{N}$ of $\mu^{N}$
This implies (4.35). From Lemma 4.8, we obtain (4.36). Indeed,
$\lim_{\mathfrak{n}\to\infty}\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\mathsf{d}_{R}^{N_{\mathfrak{n}}}d\mu_{R}^{N_{\mathfrak{n}},[1]}=\lim_{\mathfrak{n}\to\infty}\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\mathsf{d}_{R}^{N_{\mathfrak{n}}}\mathfrak{m}_{R}^{N_{\mathfrak{n}},[1]}d\Lambda_{R}^{[1]}$
$=\lim_{\mathfrak{n}\to\infty}\int_{\mathbb{R}^{d}\times\mathsf{S}}(\hat{\mathsf{f}}\otimes\mathsf{g})\mathsf{f}_{Q}^{1/2}\mathsf{d}_{R}^{N_{\mathfrak{n}}}(\mathfrak{m}_{R}^{N_{\mathfrak{n}},[1]})^{1/2}(\mathsf{f}_{Q}\mathfrak{m}_{R}^{N_{\mathfrak{n}},[1]})^{1/2}d\Lambda_{R}^{[1]}$
@@
$=\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\mathsf{d}_{R}\mathfrak{m}_{R}^{[1]}d\Lambda_{R}^{[1]}=\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\mathsf{d}_{R}d\mu_{R}^{[1]}\quad\text{ by }\mathsf{h}=(\mathsf{f}_{Q}\hat{\mathsf{f}})\otimes\mathsf{g}.$
This implies (4.36). From (4.36), (4.11), and ((A2)), we obtain
@@
$=$ $-\int_{\mathbb{R}^{d}\times\mathsf{S}}\nabla_{{x}}\mathsf{h}d\mu_{R}^{[1]}\quad\text{ by $(${A2}$)$}.$ (4.38)
Here, for the last line, we have used the weak convergence of $\mu^{N_{\mathfrak{n}}}$ to $\mu$, which imposed by ((A2)) ii. From (4.38), we deduce (4.37). Thus, we have obtained i.
We easily extend (4.37) to all $\mathsf{h}=\hat{\mathsf{f}}\otimes\mathsf{g}$ such that $\hat{\mathsf{f}}\in C_{0}^{\infty}(S_{R})$. Hence, $\mathsf{d}_{R}$ is the logarithmic derivative of $\mu_{R}$. This implies ii. ∎
We now construct the logarithmic derivative $\mathsf{d}^{\mu}$ of $\mu$ using the consistency of $\mathsf{d}_{R}$.
@@
The logarithmic derivative $\mathsf{d}^{\mu}$ of $\mu$ exists. For $\mu^{[1]}$-a.e. and in $L_{\mathrm{loc}}^{2}(\mu^{[1]})$,
$\mathsf{d}^{\mu}({x},\mathsf{s})=\lim_{R\to\infty}\mathsf{d}_{R}({x},\mathsf{s}).$ (4.39)
In particular, $\mathsf{d}^{\mu}\in L_{\mathrm{loc}}^{2}(\mu^{[1]})$ and
$\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\mathsf{d}^{\mu}d\mu^{[1]}=-\int_{\mathbb{R}^{d}\times\mathsf{S}}\nabla_{{x}}\mathsf{h}d\mu^{[1]}\quad\text{ for each $\mathsf{h}\in\mathscr{D}_{\circ\mathrm{b}}$},$ (4.40)
$E^{\mu_{{x}}}[\mathsf{d}^{\mu}({x},\mathsf{s})|\mathscr{F}_{R}]=\mathsf{d}_{R}({x},\mathsf{s})\quad\text{ for }{x}\in\overline{S}_{R},$ (4.41)
where $\mathscr{F}_{R}=\sigma[\pi_{\overline{S}_{R}}]$ and $\mu_{x}$ is the reduced Palm measure of $\mu$ conditioned at x{x}.

### v4

@@ 8.1. The $m$-labeled process for the lower Dirichlet form
$\lim_{a\to\infty}\liminf_{R\to\infty}P(\max_{1\leq i\leq m}\max_{0\leq t\leq T}\lvert\ \underline{X}_{R}^{i}(t)\rvert\leq a)=1,\quad m,T\in\mathbb{N},$ (8.12)
$\lim_{l\to\infty}\inf_{R\in\mathbb{N}}P(I_{R,T}(\underline{\mathbf{X}}_{R})\leq l)=1,\quad T\in\mathbb{N}.$ (8.13)
Here, $c_{8.1}>0$ is a constant.
Proof.
From (2.44) and (8.4), we can assume $\varphi=1$ without loss of generality. We have $\langle\int_{0}^{t}\sigma(\underline{X}_{R}^{i}(u))dB_{u}^{i}\rangle_{t}=\int_{0}^{t}\sigma^{t}\sigma(\underline{X}_{R}^{i}(u))du$ and $\sigma^{t}\sigma=\mathfrak{a}$. From (2.31), $\mathfrak{a}$ is uniformly elliptic and bounded. Hence, from (8.6) and the martingale inequality, we have (8.11)–(8.13). The proof of (8.13) is not as easy as other claims. We can prove (8.13) in the same fashion as Lemma 8.7 in [34, pp. 1212–1214] and the detail is omitted here. ∎
For a fully labeled process $\underline{\mathbf{X}}=(\underline{X}^{i})_{i\in\mathbb{N}}$, let $\underline{\mathbf{X}}^{[m]}=(\underline{\mathbf{X}}^{m},\underline{\mathsf{X}}^{m*})$ be the $m$-labeled process defined by $\underline{\mathbf{X}}^{m}=(\underline{X}^{i})_{i=1}^{m}$ and $\underline{\mathsf{X}}_{t}^{m*}=\sum_{i>m}^{\infty}\delta_{\underline{X}_{t}^{i}}$. Let $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{lwr}}^{[m]})$ be the Dirichlet form as in Lemma 7.7.
Theorem 8.1.
There exists an $\underline{\mathbf{X}}$ satisfying the following:
i For each $m\in\{0\}\cup\mathbb{N}$, $\underline{\mathbf{X}}^{[m]}$ is the $\mu^{[m]}$-symmetric, conservative, continuous Markov process associated with $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{lwr}}^{[m]})$ on $L^{2}(\mu^{[m]})$.
ii For each $m\in\{0\}\cup\mathbb{N}$, $\underline{\mathbf{X}}^{[m]}$ satisfies
$\underline{\mathbf{X}}^{[m]}=\lim_{R\to\infty}\underline{\mathbf{X}}_{R}^{[m]}\quad\text{ in law in $W^{[m]}$}.$ (8.14)
iii $\underline{\mathbf{X}}$ satisfies
$\underline{\mathbf{X}}=\lim_{R\to\infty}\underline{\mathbf{X}}_{R}\quad\text{ in law in $W(\mathbb{R}^{d})^{\mathbb{N}}$}.$ (8.15)
Proof.
@@
Let $\underline{\mathbf{X}}$ be any limit point of $\underline{\mathbf{X}}_{R}$. According to Lemma 7.4 i, $\underline{\mathbf{X}}_{R}^{[m]}$ converges to $\underline{\mathbf{X}}^{[m]}$ in law in $W^{[m]}$ and $\underline{\mathbf{X}}^{[m]}$ is a $\mu^{[m]}$-symmetric and continuous Markov process associated with $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{lwr}}^{[m]})$ on $L^{2}(\mu^{[m]})$. In particular, $\underline{\mathbf{X}}^{[m]}$ are unique in law and the convergence in (8.14) holds. Because $\underline{\mathbf{X}}_{R}^{[m]}=(\underline{\mathbf{X}}_{R}^{m},\underline{\mathsf{X}}_{R}^{m*})$ is tight in $W^{[m]}$, the limit continuous Markov process $\underline{\mathbf{X}}^{[m]}$ in (8.14) is conservative.
Because (8.14) holds for all $m\in\mathbb{N}$, the convergence in (8.15) holds and $\underline{\mathbf{X}}$ is unique in law. The second statement of iii follows from the first since $\underline{\mathbf{X}}$ is a $W^{\mathbb{N}}$-valued random variable by (8.15). ∎
8.2. Weak solutions of ISDEs for the lower Dirichlet form
Let $\underline{\mathbf{X}}$ be as in Theorem 8.1. Subsection 8.2 aims to prove that $\underline{\mathbf{X}}$ is a weak solution of ISDE (2.33) satisfying ((SIN)), ((NBJ)), and ((AC))μ.
Let $\underline{\mathbf{X}}_{R}$ be as in Lemma 8.1. Let
$\underline{\mathscr{X}}_{R}^{i}(u)=(\underline{X}_{R}^{i}(u),\sum_{j\neq i}^{\infty}\delta_{\underline{X}_{R}^{j}(u)}),\quad\underline{\mathscr{X}}_{u}^{i}=(\underline{X}^{i}_{u},\sum_{j\neq i}^{\infty}\delta_{\underline{X}_{u}^{j}}).$
Let $\mathfrak{a}\in C_{b}^{2}(\mathbb{R}^{d})$ and $\mathfrak{b}=\frac{1}{2}\{\mathrm{div}\mathfrak{a}+\mathsf{d}^{\mu}\mathfrak{a}\}$ be as in (2.31) and (2.32). Let $\sigma\in C_{b}^{1}(\mathbb{R}^{d})$ be a matrix-valued function such that $\sigma^{t}\sigma=\mathfrak{a}$,
Lemma 8.3.
The following convergence holds in law in $W(\mathbb{R}^{d})^{\mathbb{N}}$.
$\lim_{R\to\infty}\Big(\int_{0}^{\cdot}\sigma(\underline{X}_{R}^{i}(u))\,dB_{u}^{i}\Big)_{i\in\mathbb{N}}$ $=\Big(\int_{0}^{\cdot}\sigma(\underline{X}_{u}^{i})\,dB_{u}^{i}\Big)_{i\in\mathbb{N}},$ (8.16)
$\lim_{R\to\infty}\Big(\int_{0}^{\cdot}\mathfrak{b}(\underline{\mathscr{X}}_{R}^{i}(u))\,du\Big)_{i\in\mathbb{N}}$ $=\Big(\int_{0}^{\cdot}\mathfrak{b}(\underline{\mathscr{X}}_{u}^{i})\,du\Big)_{i\in\mathbb{N}}.$ (8.17)
Proof.
Let $\langle M\rangle_{t}$ be the quadratic variation process of a continuous martingale $M_{t}$. By (8.15), $\sigma^{t}\sigma=\mathfrak{a}$, and $\mathfrak{a}\in C_{b}^{2}(\mathbb{R}^{d})$, we obtain, for all $0\leq t<\infty$,
$\lim_{R\to\infty}\Big(\big\langle\int_{0}^{\cdot}\sigma(\underline{X}_{R}^{i}(u))\,dB_{u}^{i}\big\rangle_{t}\Big)_{i\in\mathbb{N}}=\lim_{R\to\infty}\Big(\int_{0}^{\cdot}\mathfrak{a}(\underline{X}_{R}^{i}(u))\,du\Big)_{i\in\mathbb{N}}$
$\quad\quad\quad\quad\quad\quad=\Big(\int_{0}^{\cdot}\mathfrak{a}(\underline{X}_{u}^{i})\,du\Big)_{i\in\mathbb{N}}=\Big(\big\langle\int_{0}^{\cdot}\sigma(\underline{X}_{u}^{i})\,dB_{u}^{i}\big\rangle_{t}\Big)_{i\in\mathbb{N}}.$ (8.18)
From (8.18), we obtain (8.16), because convergence in law in $W(\mathbb{R}^{d})$ of continuous martingales follows from that of their quadratic variations.
From Theorem 2.1, we have $\mathsf{d}^{\mu}\in L_{\mathrm{loc}}^{2}(\mu^{[1]})$. Then $\mathfrak{b}_{Q}:=1_{\overline{S}_{Q}}\mathfrak{b}\in L^{2}(\mu^{[1]})$ for any $Q\in\mathbb{N}$. Hence for any $\epsilon>0$, there exists $\mathfrak{b}_{Q}^{\epsilon}\in C_{b}(\mathbb{R}^{d}\times\mathsf{S})$ such that $\mathfrak{b}_{Q}^{\epsilon}=1_{\overline{S}_{Q}}\mathfrak{b}_{Q}^{\epsilon}$ and
Hence for any $T,i\in\mathbb{N}$,
$E\big[\int_{0}^{T}\big|(\mathfrak{b}_{Q}-\mathfrak{b}_{Q}^{\epsilon})(\underline{\mathscr{X}}_{R}^{i}(u))\big|du\big]$ $\leq E\big[\int_{0}^{T}\sum_{i\in\mathbb{N}}\big|(\mathfrak{b}_{Q}-\mathfrak{b}_{Q}^{\epsilon})(\underline{\mathscr{X}}_{R}^{i}(u))\big|du\big]$
@@ 8.2. Weak solutions of ISDEs for the lower Dirichlet form
Here we used (2.44) and (8.4), the fact that $\underline{\mathbf{X}}_{R}^{[1]}$ is a $\mu^{[1]}$-symmetric Markov process, and (8.19). Similarly, we have, for any $T,i\in\mathbb{N}$,
@@
Note that $\mathfrak{b}_{Q}^{\epsilon}\in C_{b}(\mathbb{R}^{d}\times\mathsf{S})$. For any $F\in C_{b}(W(\mathbb{R}^{d}))$, we have
Hence, from (8.15) and (8.22), we obtain, for each $i\in\mathbb{N}$,
$\lim_{R\to\infty}E\Big[F\Big(\int_{0}^{\cdot}\mathfrak{b}_{Q}^{\epsilon}(\underline{\mathscr{X}}_{R}^{i}(u))du\Big)\Big]=E\Big[F\Big(\int_{0}^{\cdot}\mathfrak{b}_{Q}^{\epsilon}(\underline{\mathscr{X}}_{u}^{i})du\Big)\Big].$
From this, (8.20), and (8.21), we easily obtain (8.17). ∎
Lemma 8.4.
For each $i,T\in\mathbb{N}$,
Proof.
From (8.15), $\{\underline{X}_{Q}^{i}(\cdot)-\underline{X}_{Q}^{i}(0)\}_{Q\in\mathbb{N}}$ is tight in $W(\mathbb{R}^{d})$. Hence,
By Lemma 8.3, $\{\int_{0}^{\cdot}\sigma(\underline{X}_{Q}^{i}(u))dB_{u}^{i}\}_{Q\in\mathbb{N}}$ and $\{\int_{0}^{\cdot}\mathfrak{b}(\underline{X}_{Q}^{i}(u),\underline{\mathsf{X}}_{Q}^{i\diamond}(u))du\}_{Q\in\mathbb{N}}$ are tight in $W(\mathbb{R}^{d})$. Similarly as (8.24), we obtain
$\lim_{R\to\infty}P\Big(\max_{0\leq t\leq T}\lvert\int_{0}^{t}\mathfrak{b}(\underline{\mathscr{X}}_{R}^{i}(u))du\rvert\geq\frac{R}{3}\Big)=0.$ (8.25)
$-\int_{0}^{t}\mathfrak{b}(\underline{\mathscr{X}}_{R}^{i}(u))du.$ (8.26)
$+P\Big(\max_{0\leq t\leq T}\lvert\int_{0}^{t}\sigma(\underline{X}_{R}^{i}(u))dB_{u}^{i}du\rvert\geq\frac{R}{3}\Big)+P\Big(\max_{0\leq t\leq T}\lvert\int_{0}^{t}\mathfrak{b}(\underline{\mathscr{X}}_{R}^{i}(u))du\rvert\geq\frac{R}{3}\Big).$
Taking $R\to\infty$ and using (8.25), we obtain (8.23). ∎
$\mathbb{X}_{R}^{i}$ $=\Big(\underline{X}_{R}^{i}(\cdot)-\underline{X}_{R}^{i}(0),\int_{0}^{\cdot}\sigma(\underline{X}_{R}^{i}(u))dB_{u}^{i},\int_{0}^{\cdot}\mathfrak{b}(\underline{\mathscr{X}}_{R}^{i}(u))du,\int_{0}^{\cdot}\frac{1}{2}\mathbf{n}_{R}\mathfrak{a}(\underline{X}_{R}^{i}(u))dL_{R}^{i}(u)\Big),$
$\mathbb{X}^{i}$ $=\Big(\underline{X}_{\cdot}^{i}-\underline{X}_{0}^{i},\int_{0}^{\cdot}\sigma(\underline{X}_{u}^{i})dB_{u}^{i},\int_{0}^{\cdot}\mathfrak{b}(\underline{\mathscr{X}}_{u}^{i})du,\mathbf{0}\Big).$ (8.27)
Lemma 8.5.
We obtain
$(\mathbb{X}^{i})_{i\in\mathbb{N}}=\lim_{R\to\infty}$ $(\mathbb{X}_{R}^{i})_{i\in\mathbb{N}}\quad\text{ in law in $C([0,\infty);(\mathbb{R}^{d})^{4})^{\mathbb{N}}$}.$ (8.28)
Proof.
$(\mathbb{X}^{i})_{i\in\mathbb{N}}=(\underline{\mathbf{X}}_{\cdot}-\underline{\mathbf{X}}_{0},\Big(\int_{0}^{\cdot}\sigma(\underline{X}_{u}^{i})dB_{u}^{i}\Big)_{i\in\mathbb{N}},\Big(\int_{0}^{\cdot}\mathfrak{b}(\underline{\mathscr{X}}_{u}^{i})du\Big)_{i\in\mathbb{N}},(\mathbf{0})_{i\in\mathbb{N}})$
We rewrite $(\mathbb{X}_{R}^{i})_{i\in\mathbb{N}}$ similarly. The convergence of $\underline{\mathbf{X}}_{R}(\cdot)-\underline{\mathbf{X}}_{R}(0)$ to $\underline{\mathbf{X}}(\cdot)-\underline{\mathbf{X}}(0)$ follows from (8.15) and (2.45). The convergence of the other terms follows from Lemmas 8.3 and 8.4. From these, we easily obtain (8.28). ∎
Theorem 8.2.
$\underline{\mathbf{X}}$ is a weak solution of (2.33) satisfying $\underline{\mathbf{X}}(0)\mathop{\sim}\limits_{\mathrm{law}}(\varphi\mu)\circ\mathfrak{l}^{-1}$, ((SIN)), ((NBJ)), and ((AC))μ.
Proof.
Let $\{h_{k}\}$ be an increasing sequence of positive numbers such that $\lim_{k}h_{k}=\infty$. Let $S_{h_{k}}=\{s\in\mathbb{R}^{d};\lvert s\rvert<h_{k}\}$. For $\mathbf{w}=(w_{i})_{i=1}^{4}\in W(\mathbb{R}^{d})^{4}$, let
$\tau_{k}(\mathbf{w})=\inf\{t>0;\mathbf{w}(t)\notin S_{h_{k}}^{4}\}.$ (8.29)
We can and do take $\{h_{k}\}$ such that, for all $i,k\in\mathbb{N}$,
$P(\mathbb{X}^{i}\in\mathbf{V}_{k})=1.$ (8.30)
$F_{k}(\mathbb{X}^{i}(t\wedge\tau_{k}(\mathbb{X}^{i})))=\lim_{R\to\infty}F_{k}(\mathbb{X}_{R}^{i}(t\wedge\tau_{k}(\mathbb{X}_{R}^{i})))\ \text{ in law}.$ (8.31)
Recall that $\underline{\mathbf{X}}_{R}=(\underline{X}_{R}^{i})_{i=1}^{\infty}$ satisfies SDE (8.1). By the definition of $F_{k}$, (8.27), (8.29), and (8.1), we obtain, for each $R,i,k\in\mathbb{N}$ and $t\in[0,\infty)$,
$\quad-\int_{0}^{t\wedge\tau_{k}(\mathbb{X}_{R}^{i})}\mathfrak{b}(\underline{\mathscr{X}}_{R}^{i}(u))du-\int_{0}^{t\wedge\tau_{k}(\mathbb{X}_{R}^{i})}\frac{1}{2}\mathbf{n}_{R}\mathfrak{a}(\underline{X}_{R}^{i}(u))dL_{R}^{i}(u)=0.$
From this and (8.31), we obtain, for each $i,k\in\mathbb{N}$ and $t\in[0,\infty)$,
$F_{k}(\mathbb{X}^{i}(t\wedge\tau_{k}(\mathbb{X}^{i})))=0,$
which implies that
$\underline{X}_{t\wedge\tau_{k}(\mathbb{X}^{i})}^{i}-\underline{X}_{0}^{i}-\int_{0}^{t\wedge\tau_{k}(\mathbb{X}^{i})}\sigma(\underline{X}_{u}^{i})dB_{u}^{i}-\int_{0}^{t\wedge\tau_{k}(\mathbb{X}^{i})}\mathfrak{b}(\underline{\mathscr{X}}_{u}^{i})du=0.$
By (8.27) and Lemma 8.5, the equation above holds as the identity of continuous processes. Hence, using $\lim_{k\to\infty}\tau_{k}(\mathbb{X}^{i})=\infty$ yields ISDE (2.33).
To prove the non-collision property of $\underline{\mathbf{X}}$, we use Proposition 6.2. From Theorem 2.1, $\mathsf{d}^{\mu}\in L_{\mathrm{loc}}^{2}(\mu^{[1]})$. From Theorem 8.1, the one-labeled process $\underline{\mathbf{X}}^{[1]}$ given by $\underline{\mathbf{X}}$ is a $\mu^{[1]}$-symmetric Markov process. Hence, the $\mu^{[1]}$-symmetry of $\underline{\mathbf{X}}^{[1]}$ and (2.44) and (8.4) imply that, for all $R,T\in\mathbb{N}$,
@@
From this, we obtain (6.1). Note that (6.14) follows from (2.44) and (8.4).
Thus, we have verified the assumptions (6.1) and (6.14) of Proposition 6.2. Hence, we obtain the non-collision property (6.3). Thus, we obtain ((SIN)).
From (8.6) and Theorem 8.1, we have, for each $i\in\mathbb{N}$,
$\underline{X}_{t}^{i}-\underline{X}_{0}^{i}=\frac{1}{2}\Big\{\int_{0}^{t}\sigma(\underline{X}_{u}^{i})dB_{u}^{i}-\int_{0}^{t}\sigma(\underline{X}_{u}^{i})dB_{u}^{i}\circ r_{t}\Big\}.$ (8.32)
Assumption ((AC))μ is obvious because $\underline{\mathsf{X}}$ is associated with the reversible Dirichlet form $(\mathscr{E}^{\mu},\underline{\mathscr{D}}^{\mu})$ on $L^{2}(\mu)$ and $\underline{\mathsf{X}}_{0}\mathop{\sim}\limits_{\mathrm{law}}\varphi\mu$. ∎
The aim of Section 9 is to prove Theorem 2.2. In Section 9.1, we construct the $m$-labeled diffusion $\overline{\mathbf{X}}^{[m]}$. In Section 9.2, we then construct the associated fully labeled process $\overline{\mathbf{X}}$ and show that $\overline{\mathbf{X}}$ is a weak solution of (2.33) satisfying ((SIN)), ((NBJ)), and ((AC))μ. These conditions are essential for establishing pathwise uniqueness and the existence of strong solutions to the ISDE (2.33).
Unlike in Section 8, the upper Dirichlet form $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}}^{[m]})$ is quasi-regular. Therefore, we construct the $m$-labeled diffusion $\overline{\mathbf{X}}^{[m]}$ directly from the infinite-volume upper Dirichlet form. This approach differs markedly from the finite-volume approximation used in Section 8.
All results in Section 9 are proved under Assumptions ((A1))((A3)).
9.1. The $m$-labeled process for the upper Dirichlet form
Let $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}}^{[m]})$ be the Dirichlet form on $L^{2}(\mu^{[m]})$ as in Lemma 7.7. Let $\mathrm{Cap}^{[m]}$ denote the capacity with respect to $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}}^{[m]})$ on $L^{2}(\mu^{[m]})$. We write $\mathrm{Cap}^{\mu}$ for the case $m=0$. Define $\mathfrak{u}^{[m,0]}\!:\!(\mathbb{R}^{d})^{m}\times\mathsf{S}\!\to\!\mathsf{S}$ by $\mathfrak{u}^{[m,0]}(\mathbf{x},\mathsf{s})=\mathfrak{u}(\mathbf{x})+\mathsf{s}$.
Lemma 9.1.
Let $A\subset\mathsf{S}$. If $\mathrm{Cap}^{\mu}(A)=0$, then, for each $m\in\mathbb{N}$,
$\mathrm{Cap}^{[m]}\bigl((\mathfrak{u}^{[m,0]})^{-1}(A)\bigr)=0.$
Proof.
The proof follows from the argument of [26, Lem. 4.1]. ∎
@@ 9.1. The $m$-labeled process for the upper Dirichlet form
i There exists a $\mu^{[m]}$-symmetric, conservative diffusion $\overline{\mathbf{X}}^{[m]}$ properly associated with $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}}^{[m]})$ on $L^{2}(\mu^{[m]})$ for each $m\in\{0\}\cup\mathbb{N}$.
ii None of the labeled particles $\overline{X}^{i}$, where $i=1,\ldots,m$, in $\overline{\mathbf{X}}^{[m]}$ explode or collide with other labeled particles:
$P(\overline{X}_{t}^{i}\neq\overline{X}_{t}^{j}\text{ for all }t\in[0,\infty),1\leq i,j\leq m,i\neq j)=1.$ (9.2)
iii For each $1\leq i\leq m$,
$\overline{X}_{t}^{i}-\overline{X}_{0}^{i}=\int_{0}^{t}\sigma(\overline{X}_{u}^{i})\,dB_{u}^{i}+\int_{0}^{t}\mathfrak{b}(\overline{\mathscr{X}}_{u}^{i})\,du,$ (9.3)
$\overline{X}_{t}^{i}-\overline{X}_{0}^{i}=\frac{1}{2}\Big\{\int_{0}^{t}\sigma(\overline{X}_{u}^{i})\,dB_{u}^{i}-\int_{0}^{t}\sigma(\overline{X}_{u}^{i})\,dB_{u}^{i}\circ r_{t}\Big\}.$ (9.4)
Here $r_{t}$ is the time-reversal operator defined before Lemma 8.1, and
$\overline{\mathscr{X}}_{u}^{i}=\big(\overline{X}_{u}^{i},\sum_{j\neq i}^{\infty}\delta_{\overline{X}_{u}^{j}}\big).$ (9.5)